# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

### v1

@@ B.2 CR Metric
B.3 MIACR Metric
Table 15 presents the performance of seven machine unlearning methods on CIFAR-10 in ResNet-18, evaluated with the MIACR metric. In addition to the settings discussed in Section 4, we include results for $\alpha\in[0.05,0.15]$ in Table 15.
@@ B.3 MIACR Metric
B.4 How can we better measure forgetting under distribution shifts
Recall that the coverage value of RL has a relatively large deviation in Table 5 since it employs label corruption in its unlearning strategy which can cause distribution shifts. Here, we introduce how to better measure forgetting under these circumstances.
Figure 2(a) shows the non-conformity score distribution of calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ in unlearning model $\theta_{u}$ obtained by RL method in Tiny ImageNet with ViT. It looks like there is a significant discrepancy between the distribution of the forget data and the calibration data.
@@ Appendix C Performance of Our Unlearning Framework
Furthermore, we conduct an ablation study and analyze the impact of using our unlearning framework. As illustrated in Figures 3 and 4, we compare the density distributions of non-conformity scores for calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ under the RT, FT, and RL unlearning methods. We set $\lambda$ to 1. Clearly, a higher non-conformity score for $\mathcal{D}_{f}$ indicates that it is less likely to be included in the conformal prediction set, reflecting more effective forgetting. Comparing Figures 3 and 4, it is evident that after applying our unlearning framework, the distribution of non-conformity scores for forget data shifts noticeably to the right. This demonstrates the effectiveness of our framework in enhancing forgetting quality.
Table 9: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario. The results are reported in the format a±$\pm$±b, where a is the mean and b is the standard deviation from $3$ independent trials. The performance gap relative to Retrain method is represented in ().
@@
FF 0.05 $0.973_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $7.966_{\pm 0.212}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.877})$ $7.408_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.334})$ $0.122_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.742})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.751})$ $0.999_{\pm 0.000}$
0.1 $0.933_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129_{\pm 0.148}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566_{\pm 0.166}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998_{\pm 0.001}$
UA59.9%, RA40.1%, TA41.1% 0.15 $0.888_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.068})$ $0.852_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $6.431_{\pm 0.078}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.590})$ $5.903_{\pm 0.262}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.036})$ $0.138_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.837})$ $0.144_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.824})$ $0.996_{\pm 0.002}$
0.2 $0.835_{\pm 0.048}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.794_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $5.750_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.961})$ $5.219_{\pm 0.368}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.395})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.843})$ $0.153_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.828})$ $0.993_{\pm 0.003}$
@@
Table 10: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in $50\%$ random data forgetting scenario.
@@
FF $0.05$ $0.972_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.954_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $8.023_{\pm 0.189}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.736})$ $7.461_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.247})$ $0.121_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.621})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.652})$ $0.999_{\pm 0.000}$
$0.1$ $0.930_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $7.091_{\pm 0.151}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.068})$ $6.521_{\pm 0.171}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.501})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.747})$ $0.138_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.748})$ $0.998_{\pm 0.001}$
UA60.0%, RA40.1%, TA40.6% $0.15$ $0.887_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.852_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $6.402_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.519})$ $5.837_{\pm 0.438}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.931})$ $0.139_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.146_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.788})$ $0.996_{\pm 0.001}$
$0.2$ $0.840_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.803_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $5.805_{\pm 0.042}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.994})$ $5.253_{\pm 0.376}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.403})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.819})$ $0.153_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.993_{\pm 0.003}$
@@
Table 11: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in class-wise forgetting scenario.
@@
FF 0.05 $0.984_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.962_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.950_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $7.760_{\pm 1.386}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.240})$ $7.479_{\pm 1.223}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.521})$ $7.627_{\pm 0.058}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.479})$ $0.129_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.130_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.125_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.715})$ $0.999_{\pm 0.000}$ $0.997_{\pm 0.003}$
0.1 $0.964_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.917_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.909_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $7.360_{\pm 0.999}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.640})$ $6.988_{\pm 0.749}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.012})$ $6.853_{\pm 0.390}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.930})$ $0.132_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.132_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.133_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.823})$ $0.998_{\pm 0.000}$ $0.996_{\pm 0.003}$
UA100%, UAtf100%, 0.15 $0.922_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.852_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.863_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $6.941_{\pm 1.088}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.059})$ $6.562_{\pm 0.856}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.438})$ $6.278_{\pm 0.232}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.395})$ $0.134_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.131_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.138_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.997_{\pm 0.001}$ $0.994_{\pm 0.006}$
RA34.3%, TA36.1% 0.2 $0.897_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.818_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.182})$ $0.789_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $6.757_{\pm 0.978}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.243})$ $6.352_{\pm 0.757}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.648})$ $5.307_{\pm 0.527}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.477})$ $0.134_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.130_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.150_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.993_{\pm 0.002}$ $0.993_{\pm 0.006}$

### v2

@@ B.3 CR Metric
B.4 MIACR Metric
Table 13 presents the performance of 9 machine unlearning methods on CIFAR-10 in ResNet-18, evaluated with the MIACR metric. In addition to the settings discussed in Section 4, we include results for $\alpha\in[0.05,0.15]$ in Table 13.
@@ B.4 MIACR Metric
B.5 Measuring Forgetting under Distribution Shifts
RL and Salun are unlearning methods employ label corruption in their unlearning strategy which can cause distribution shifts. Here, we introduce how to better measure forgetting under these circumstances. Figure 4(a) shows the non-conformity score distribution of calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ in unlearning model $\theta_{u}$ obtained by RL method in Tiny ImageNet with ViT. It looks like there is a significant discrepancy between the distribution of the forget data and the calibration data.
@@ Appendix C Performance of Our Unlearning Framework
Furthermore, we conduct an ablation study and analyze the impact of using our unlearning framework. As illustrated in Figures 5 and 6, we compare the density distributions of non-conformity scores for calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ under the RT, FT, and RL unlearning methods. We set $\lambda$ to 1. Clearly, a higher non-conformity score for $\mathcal{D}_{f}$ indicates that it is less likely to be included in the conformal prediction set, reflecting more effective forgetting.
Comparing Figures 5 and 6, after applying our unlearning framework, we observe a significant rightward shift in the non-conformity score distribution of forget data, which is a promising signal according to evaluation criterion ➋. Furthermore, the FT distribution in Figure 6 exhibits substantial overlap with the calibration data, nearly matching the distribution observed in RT. Based on evaluation criterion ➊, since calibration data represents unseen examples, the similarity between forget data and calibration data distributions provides strong evidence of effective forgetting. Overall, the results evaluated on both evaluation criteria ➊ and ➋ consistently confirm the efficacy of our framework in enhancing forgetting quality.
Appendix D Limitations
While our work provides a new perspective on evaluating and improving machine unlearning via uncertainty quantification, there are a few limitations to consider.
First, our current study focuses exclusively on image classification tasks. While this setting is representative and widely used in prior unlearning literature Foster et al. (2024); Graves et al. (2021); Shen et al. (2024), extending our framework to other domains such as language modeling or generative models remains an important direction for future work.
Second, our evaluation method requires a separate validation set for uncertainty calibration. But this requirement is generally not restrictive in real-world unlearning settings, as model controllers (e.g., service providers) typically have full access to the training and validation data. Moreover, in the scarce training set scenario, where access to data is inherently limited, evaluation is often unnecessary because retraining from scratch is computationally feasible and serves as a natural baseline. Thus, the need for calibration data does not pose a significant barrier to applying our approach in most practical contexts.
Appendix E Impact Statement
Our work on machine unlearning through uncertainty quantification perspective offers several positive contributions to society.
Enhanced Privacy Protection.
Effective machine unlearning is crucial for upholding individual privacy rights, particularly in contexts where regulations like GDPR mandate the "right to be forgotten." Our method provides a mechanism to verify that unlearning is genuine, ensuring that user data is truly removed from models when requested.
Support for Trustworthy AI Development.
Our approach provides technical foundations for AI systems capable of "forgetting" harmful, biased, or outdated information, facilitating continuous improvement and the trustworthiness of AI systems.
(a) FT
(b) RL
(c) Salun
(d) SFRon
[…]
(a) FT
(b) RL
(c) Salun
(d) SFRon
[…]
(a) FT
(b) RL
(c) Salun
(d) SFRon
[…]
(a) FT
(b) RL
(c) Salun
(d) SFRon
[…]
Table 7: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario. The results are reported in the format a±$\pm$±b, where a is the mean and b is the standard deviation from $3$ independent trials. The performance gap relative to Retrain method is represented in ().
@@ Appendix E Impact Statement
SFRon 0.05 $0.977_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.953_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.100_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $1.143_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.889_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.834_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.926_{\pm 0.018}$
0.1 $0.945_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.905_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.986_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.977_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.958_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.927_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.435_{\pm 0.043}$
UA4.8%, RA97.4%, TA91.4% 0.15 $0.895_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.847_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.912_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.879_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.982_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.963_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.082_{\pm 0.007}$
0.2 $0.857_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.808_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.868_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $0.826_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.988_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.978_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.025_{\pm 0.005}$
Table 8: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet18 in $50\%$ random data forgetting scenario.
@@
SFRon 0.05 $0.977_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.953_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.100_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.188})$ $1.143_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.889_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.147})$ $0.834_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.926_{\pm 0.018}$
0.1 $0.945_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.905_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.986_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.977_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.044})$ $0.958_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $0.927_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.435_{\pm 0.043}$
UA4.0%, RA97.3%, TA91.6% 0.15 $0.895_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.847_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.912_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.879_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.982_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.963_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.082_{\pm 0.007}$
0.2 $0.857_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.808_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.868_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.826_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.988_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.978_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.025_{\pm 0.005}$
Table 9: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet18 in class-wise forgetting scenario.
@@
NegGrad+ 0.05 $0.989_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.961_{\pm 0.056}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.945_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $9.432_{\pm 0.803}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.568})$ $9.038_{\pm 1.360}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.962})$ $1.053_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.105_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.107_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.897_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $1.000_{\pm 0.000}$ $0.835_{\pm 0.085}$
0.1 $0.980_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.954_{\pm 0.065}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.881_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $9.250_{\pm 1.061}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.750})$ $8.836_{\pm 1.647}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.164})$ $0.913_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.106_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.109_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.965_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.000_{\pm 0.000}$ $0.057_{\pm 0.021}$
UA96.2%, UAtf95.2%, 0.15 $0.952_{\pm 0.068}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.908_{\pm 0.130}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.849_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $8.600_{\pm 1.980}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.400})$ $8.077_{\pm 2.719}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.923})$ $0.868_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.113_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.116_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.977_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.000_{\pm 0.000}$ $0.012_{\pm 0.003}$
RA97.6%, TA92.8% 0.2 $0.958_{\pm 0.060}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.921_{\pm 0.111}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $0.814_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $8.673_{\pm 1.876}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.327})$ $8.219_{\pm 2.519}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.781})$ $0.828_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.112_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.115_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.983_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.000_{\pm 0.000}$ $0.004_{\pm 0.003}$